a golf ball with a mass of 0.060 kg is subjected to an impulse that causes it to move with a velocity of 60m/s. If the golf club and the ball are in contact for 0.0050 s,what is the average net force that acts on the ball?


1
Expert's answer
2021-04-25T03:50:25-0400

The impulse is the change in momentum. The direction of travel of the struck ball is the positive direction.

Δp = mΔv = 0.060 kg x (60-0) = 3.6 kg x m/s

The average force is the impulse divided by the interaction time.

F = Δp / Δt = 3.6 / 0.0050 = 720 N